OpenStudy (pphalke):

the formula m=1/5t can be used to determine how far in miles, you are from lightning when its thunder takes t seconds to reach your ears. if it takes 9 secs for the sound of thunder to reach you after you have seen the lightning how far away is the storm? will fan and medal

OpenStudy (janu16):

M = 1/5 t t = 9 So M =1/5 *9 = 1.8 miles So storm is 1.8 miles away

OpenStudy (janu16):

do you understand?

OpenStudy (janu16):

@pphalke

OpenStudy (pphalke):

yes thanks this is a practice test, i take the real one monday and have one more chance left

OpenStudy (pphalke):

@Janu16

OpenStudy (janu16):

ohok

OpenStudy (pphalke):

if i blow this then i cant pre test into the next module. how can a class be an hour and 50 mins not giving enough time for a test?

OpenStudy (janu16):

lol

OpenStudy (pphalke):

4.2x-5.25=6.3-3.5x help @Janu16

OpenStudy (janu16):

simplfy this?

OpenStudy (pphalke):

solve

OpenStudy (janu16):

ok solve for x tight?

OpenStudy (pphalke):

yes

OpenStudy (janu16):

you have any chcoies?

OpenStudy (janu16):

Add 5.255.25 to both sides 4.2x=6.3−3.5x+5.25 4.2x=6.3−3.5x+5.25 2 Simplify 6.3−3.5x+5.256.3−3.5x+5.25 to −3.5x+11.55−3.5x+11.55 4.2x=−3.5x+11.55 4.2x=−3.5x+11.55 3 Add 3.5x3.5x to both sides 4.2x+3.5x=11.55 4.2x+3.5x=11.55 4 Simplify 4.2x+3.5x4.2x+3.5x to 7.7x7.7x 7.7x=11.55 7.7x=11.55 5 Divide both sides by 7.77.7 x=1.5

OpenStudy (janu16):

i got 1.5

OpenStudy (janu16):

@pphalke

OpenStudy (pphalke):

thanks. so i have more questions can u help answer plz @Janu16

OpenStudy (janu16):

i can try an djust post it on other thread cause this one is long

OpenStudy (janu16):

and tag me and i will see what i can do

OpenStudy (pphalke):

ok